OpenStudy (msmr):

A sphere and a right cylinder have equal volumes. The radius of the sphere = the radius of the cylinder. Compare the surface area of the sphere to the total area of the cylinder.

OpenStudy (unklerhaukus):

2:3

OpenStudy (msmr):

That is also what I got; but my answer key says "6:7" and I have no idea why.

OpenStudy (unklerhaukus):

ops 2:3 is the ratio of the volumes if r is constant, the question is asking for the ratio of surface areas

Directrix (directrix):

Volume of Sphere = Volume of Right Circular Cylinder 4/3* π r ³ = π r ² h 4r/3 * π r ² = π r ² * h 4r/3 = h 4r = 3h ----------- Ratio of Surface Area of Sphere to Surface Area of Right Circular Cylinder 4π r ² / (2 π r h + 2π r²) 2r * 2 π r / 2 π r * (h + r) 2r / (h + r) 3/3 * 2r / (h + r) 6r / 3 (h + r) 6r / (3h + 3r) Recall that 4r = 3h or 3h = 4r 6r / (4r + 3r) 6r / 7r 6/7 --> Ratio of Surface Areas of Sphere and Right Circular Cylinder with Equal Volumes

OpenStudy (msmr):

Unklerhaukus: I found the ratio of the volumes too, haha. Thanks though! Directix: that helped a ton :)